# Benchmarking General Purpose In-Context Learning

### v2

@@ Abstract
In-context learning (ICL) capabilities are becoming increasingly appealing for building general intelligence due to their sample efficiency and independence from artificial optimization skills. To enhance generalization, biological neural systems primarily inherit learning capabilities and subsequently refine their memory, acquiring diverse skills and knowledge through extensive lifelong experiences. This process gives rise to the concept of general-purpose in-context learning (GPICL). Compared to standard ICL, GPICL addresses a broader range of tasks, extends learning horizons, and starts at a lower zero-shot baseline. We introduce two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark includes a vast number of tasks characterized by significant task variance and minimal transferable knowledge among tasks, facilitating lifelong in-context learning through continuous generation and interaction. These features pose significant challenges for models that rely on context or interactions to improve their proficiency, including language models, decision models, and world models. Our experiments reveal that parameter scale alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds to: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: source code available at https://github.com/FutureAGI/L3C
@@ 1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pretraining and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations). However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
@@
Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation. This disparity is further highlighted by the phenomenon known as the genomics bottleneck. Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch and Schmidhuber, 2021; Wang et al., 2022). A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 2. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
@@
Why GPICL? The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models. First, inspired by the genomics bottleneck and building on previous work (Kirsch and Schmidhuber, 2021), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills. Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model. In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks. Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
Building on those points, there can be an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans. A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
@@
  • Vast number of variant tasks with minimal knowledge transfer: To minimize the effectiveness of zero-shot generalization, it is advantageous to engage with a variety of tasks where task-specific knowledge cannot be transferred among different tasks. This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than its ability of zero-shot generalization. It is essential to require numerous tasks to prevent the meta-model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that most meta-learning benchmarks fall into when given limited tasks.
@@
  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, such that in-context reinforcement learning capabilities can be effectively meta-trained and evaluated. Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
@@
  • Lifelong in-context learning: GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”). It is necessary to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”. While many previous gradient-based meta-learning methods also encode adaptation over millions of samples (e.g., many gradient-based meta-learning methods), most of these are to blame on the inefficient learning functionality. Consequently, they should not be considered as representing a long ICL horizon.
@@
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL. One crucial conclusion from these limited experiments may be that parameter scale alone is not crucial for ICL or GPICL. This finding aligns with previous research indicating that the size of contexts and memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022). This insight may highlight alternative approaches beyond merely scaling up models within the current AI community.
@@ 2.2 Baseline Solution
As the training data, we generated 500K sequences with a pre-defined vocabulary size of $N=32$, and the complexity of the pseudo-language was uniformly sampled across $n\in{3,4,5,6}$. Each sequence comprises $4,096$ tokens, totaling 2B tokens. We apply the auto-regressive transformer (Vaswani et al., 2017) with rotary embedding (Su et al., 2024). The hyper-parameter settings include: a tiny-sized transformer (303k parameters), a small-sized transformer (9.5m parameters), and a standard-sized(151m parameters) model. The evaluation set consists of independently sampled sequences, totaling 4K samples (16M tokens) for each $n\in\{2,3,4,5,6,7,8\}$, sampled uniformly. We assessed the model’s performance by plotting the averaged perplexity ($-\log p(x_{t})$) for each position across different $n$ values. To assess how the trained model generalizes, we also included the PG-19 (Rae et al., 2019) test set as a natural corpus. In this set, words are broken down into 26 alphabetic characters and randomly mapped onto 32 pre-defined tokens, the left slot are given to punctuation marks. The entire test texts are concatenated into a long sequence and broken down into sentences of 4,096 characters each. It is important to note that the model has never been trained on any natural language corpora. Specific details are provided in the appendices.
@@ 2.3.1 Towards Meta-Language Model
We demonstrate that this pseudo-language trained model can be regarded as a meta-language model designed to adapt to any conceivable language through GPICL. See from basic evaluation results of 151M model on the test sets in Figure 3, several observations are noteworthy: First, perplexity improves with increasing context length across all values of $n$, indicating that the model is effectively leveraging in-context learning. Notably, for $n=7$ and $n=8$—complexities on which the model was not explicitly trained—we observed a consistent semi-logarithmic improvement in perplexity with increasing context length. Second, a higher language complexity, indicated by larger values of $n$, correlates with a more gradual improvement in perplexity relative to context length. This phenomenon confirms that increased language complexity introduces longer-term dependencies within contexts, establishing meta language as an ideal benchmark for evaluating long-dependency modeling capabilities. Most importantly, the model demonstrates a significant performance gain through in-context learning (ICL) in realistic natural language scenarios, despite never being exposed to a single real-world natural word, illustrating that the ICL capability acquired through meta-language is not restricted to a specific language type.
(a) English vocabulary learning task.
(b) Mathematical operation task.
@@ 2.3.2 Scaling Laws for GPICL
Our experiments suggest a shift in the GPICL paradigm from “the larger the parameters, the better” to “the longer the context (and the larger the memory), the better”, provided the parameters are sufficient. In contrast, current large language models (LLMs) typically require at least billions of parameters, the majority of which we suspect are employed for zero-shot generalization, while the parameters dedicated to Incremental Context Learning (ICL) are relatively sparse. This perspective is consistent with the concept of the genomics bottleneck (Wang et al., 2022), which advocates for models that, while adequately parameterized, prioritize larger memory capacities to optimize GPICL. However, it is important to note that while the transformer architecture offers a “full memory” capability for sequence modeling, it also incurs a computational cost of $\Theta(T^{2})$. This underscores the necessity for developing more efficient transformers, which would allow for the extension of context length to longer spans. We believe research in this area (Tay et al., 2022) could benefit from this benchmark.
@@ 3.1 Problem Setting
The Maze World simulation tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route. This setting could potentially reflect a wide range of real-world tasks, such as household robots adapting to unfamiliar indoor environments. As the agent gains more experience within a specific environment, it should progressively gather more information and navigate to specified targets more efficiently. The proposed benchmarks has the following features:
@@
The agent receives a positive reward for reaching the specified target, while a minor cost is incurred for each step taken. Available actions include moving forward, moving backward, stopping, turning left, and turning right. Beyond these task settings, the benchmarks provide a wide range of hyper-parameters to generate diverse tasks. Examples include the height and angle of the perspective, ceiling height, and the length of the cell unit. In addition to command-based navigation, we also offer a “Survival” mode, details are provided in the appendices.
To optimize the initiation of model training, we have implemented rule-based smart agents, which are allowed to access the global map, but with limitations. They record everything within their line of sight into memories, while areas out of sight remain hidden. This kind of agent efficiently navigates to its goals by utilizing an exploration-then-exploitation strategy. Although this rule-based smart agent may not consistently reach the global optimum, primarily due to shortcomings in its exploration strategy, it nonetheless provides a solid baseline for imitation learning.
To model varied agent behaviors, we differentiate between Long-Term Memory (LTM) and Short-Term Memory (STM) in the rule-based agents. The STM retains only the most recent three steps of observations, while the LTM retains information until the episode concludes. We introduce a hyper-parameter, $p(\text{STM}\rightarrow\text{LTM})$, to govern the probability of transferring information from STM to LTM. At $p(\text{STM}\rightarrow\text{LTM})=0\%$, the agent operates without LTM, whereas $p(\text{STM}\rightarrow\text{LTM})=100\%$ allows for a fully functional LTM.
@@
(b) A sketch of an agent’s LTM
Figure 6: A demonstration of the maze world tasks alongside a rule-based navigation system. The rule-based agent is provided with partial access to the environment’s ground truth (b) under specific conditions. Only the areas within the agent’s line of sight are revealed, while the unsighted regions remain obscured. The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability, denoted as $p(\textit{STM}\rightarrow\textit{LTM})$ The agent employs a strategy that balances exploration with exploitation.

### v5

# Benchmarking General-Purpose In-Context Learning

@@ Abstract
In-context learning (ICL) empowers generative models to address new tasks effectively and efficiently on the fly, without relying on any artificially crafted optimization techniques. In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General-Purpose In-Context Learning (GPICL). To this end, we introduce two lightweight benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark encompasses a vast number of tasks characterized by significant task variance, facilitating meta-training that minimizes inductive bias. These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction. These characteristics necessitate the models to leverage contexts and history interactions to enhance their capabilities, across domains such as language modeling, decision-making, and world modeling. Our experiments on the baseline models demonstrate that meta-training with minimal inductive bias and ICL from the ground up is feasible across all the domains we’ve discussed. Additionally, our findings indicate that the scale of parameters alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds with: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: Source code available at https://github.com/FutureAGI/L3C
@@ 1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pre-training and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations). However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
@@
Figure 1: Conceptual diagram of in-context learning (ICL) and general-purpose in-context learning (GPICL)
There are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). The enormous lifelong learning and adaptation potential of biological neural networks, forming abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy), are also used to guide the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024). A promising area in this line of research is general-purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 1. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
@@
Importance of GPICL. The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, suppressing zero-shot capability seem to contradict the prevailing trend in large language models. To elucidate its importance further, we emphasize three key points:
First, theoretically, the lower zero-shot capability in meta-training indicates a reduced reliance on inductive biases, thereby demonstrating sufficient diversity of meta-training tasks. As a result, the meta-training process focus primarily on the functionality of ICL instead of task-specific knowledge. Conversely, the attainment of zero-shot capability might suggest that the tasks possess a common knowledge base. While it could lead to faster adaption within the scope of these tasks, it might potentially restrict the model’s applicability beyond the scope of these tasks.
Second, practically, meta-training the learning functionality requires an enormous number of tasks, which is challenging to collect in the real world. The use of abundant synthetic data in pre-training and meta-training can be essential, but tend to be non-factual and biased, presenting challenges such as the sim-to-real gap. Ideally, the meta-training stage of GPICL results in the loss of much task-specific knowledge while preserving the ICL functionality, thus helps prevent the model from the biases in the synthetic datasets. Building on this point, an alternative approach could be developed to replace the current mainstream learning pipeline of Pre-training & Fine-tuning & ICL with Meta-training & GPICL (Figure 2). In the new pipeline, initially, massive synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned, real-world data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
Third, since a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, it is challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for analysis on model architectures and learning algorithms.
@@
  • Vast number of variant tasks with minimal inductive bias: It is desirable to ensure that the meta-training and meta-testing tasks are sufficiently diverse, preventing the transfer of task-specific knowledge among different tasks. This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than inductive biases. It is also essential to require numerous tasks to prevent the model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that many meta-learning benchmarks fall into when given limited tasks, as is discussed in Kirsch et al. (2022).
@@
  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, to ensure obtaining complex functionalities such as in-context reinforcement learning. Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
@@
  • Lifelong in-context learning: To guarantee the capability of adapting complex tasks, GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”). It is important to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”, excluding scenarios requiring extensive contexts that are primarily due to inefficient learning functionalities.
@@
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone and its variants. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL. One crucial conclusion from these limited experiments may be that the scale of parameters alone is not crucial for ICL or GPICL. This finding aligns with previous research indicating that the size of contexts, and the scale of the memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022) in learning from contexts.
@@ 2.2 Baseline Solution
As the training data, we generated 500K sequences with a pre-defined vocabulary size of $N=32$, and the complexity of the pseudo-language was uniformly sampled across $n\in{3,4,5,6}$. Each sequence comprises $4,096$ tokens, totaling 2B tokens. We apply the auto-regressive transformer (Vaswani et al., 2017) with rotary embedding (Su et al., 2024). The hyper-parameter settings include: a tiny-sized transformer (303k parameters), a small-sized transformer ($9.5$M parameters), and a standard-sized(151m parameters) model. The evaluation set consists of independently sampled sequences, totaling 4K samples (16M tokens) for each $n\in\{2,3,4,5,6,7,8\}$, sampled uniformly. We assessed the model’s performance by plotting the averaged perplexity ($-\log p(x_{t})$) for each position across different $n$ values. To assess how the trained model generalizes, we also included the PG-19 (Rae et al., 2019) test set as a natural corpus. In this set, words are broken down into 26 alphabetic characters and randomly mapped onto 32 pre-defined tokens, the left slot are given to punctuation marks. The entire test texts are concatenated into a long sequence and broken down into sentences of 4,096 characters each. It is important to note that the model has never been trained on any natural language corpora. Details of the model description can be found in Appendix A.1.
@@ 2.3.1 Towards Meta-Language Model
We demonstrate that this pseudo-language trained model can be regarded as a meta-language model designed to adapt to any conceivable language through GPICL. See from basic evaluation results of 151M model on the test sets in Figure 3, several observations are noteworthy: First, perplexity improves with increasing context length across all values of $n$, indicating that the model is effectively leveraging in-context learning. Notably, for $n=7$ and $n=8$—complexities on which the model was not explicitly trained—we observed a consistent semi-logarithmic improvement in perplexity with increasing context length. Second, a higher language complexity, indicated by larger values of $n$, correlates with a more gradual improvement in perplexity relative to context length. This phenomenon confirms that increased language complexity introduces longer-term dependencies within contexts, establishing meta language as an ideal benchmark for evaluating long-dependency modeling capabilities. Most importantly, the model demonstrates a significant performance gain through in-context learning (ICL) in realistic natural language scenarios, despite never being exposed to a single real-world natural word, illustrating that the ICL capability acquired through meta-language is not restricted to a specific language type. We further test the ability of generation of the meta-language model on two rudimentary real-world language tasks, showing that in contrast to the absence of any natural language corpora, the model is able to learn to adapt to the real world language tasks through ICL only (Appendix A.3).
@@ 2.3.2 Scaling Laws for GPICL
Our experiments suggest a shift in the GPICL paradigm from increasing the parameters to increasing the context length, memory, and states provided the parameters are sufficient. In contrast, current large language models (LLMs) typically require at least billions of parameters, the majority of which we suspect are employed for zero-shot generalization, while the parameters dedicated to Incremental Context Learning (ICL) are relatively sparse. This perspective is consistent with the concept of the genomics bottleneck (Wang et al., 2022), which advocates for models that, while adequately parameterized, prioritize larger memory capacities to optimize GPICL. However, it is important to note that while the transformer architecture offers a “full memory” capability for sequence modeling, it also incurs a computational cost of $\Theta(T^{2})$. This underscores the necessity for developing more efficient transformers, which would allow for the extension of context length to longer spans. We believe research in this area (Tay et al., 2022) could benefit from this benchmark.
@@ 3.1 Problem Setting
Mazes are typical environments that can scale up to an unlimited number of layouts, features, targets, and tasks. Many previous works benchmark meta-learning with mazes, but these are limited to 2-D observation space (Wang et al., 2022; Morad et al., 2023; Nikulin et al., 2023), text-only (Ding et al., 2024) observation and action space, and limited number of layouts and tasks (Duan et al., 2016; Morad et al., 2023; Ding et al., 2024). To enhance the task diversity, we propose the Maze World benchmark. It tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route. This setting could potentially reflect a wide range of real-world tasks, such as household robots adapting to unfamiliar indoor environments. As the agent gains more experience within a specific environment, it should progressively gather more information and navigate to specified targets more efficiently. The proposed benchmarks has the following features:
@@
The agent receives a positive reward for reaching the specified target, while a minor step-cost is incurred for each step taken. Available actions include moving forward, moving backward, stopping, turning left, and turning right. Note that unlike previous works (Chen et al., 2021; Kirsch et al., 2023; Lin et al., 2023; Lee et al., 2024), we did not use rewards as inputs, nor did we explicitly predict rewards. In the “NAVIGATION” tasks of Maze World, instant rewards are easily inferred from observations. However, in other tasks within Maze World, such as “SURVIVAL” tasks (see details in the Appendix A.5), modeling rewards may be necessary. This should not impede the analysis of the benchmark and GPICL process.
To provide a cost-effective and scalable reference dataset for meta-training GPICL, we have implemented rule-based agents with a robust simulated policy. These agents are allowed to access the global maps, but with limitations: They are allowed to record map areas within their line of sight into memory, while other areas remain hidden. These agents are referred to as privileged agents because they have access to additional information. The privileged agents efficiently navigates to their goals by utilizing an exploration-then-exploitation strategy. Although these agents may not consistently reach the global optimum due to shortcomings in its exploration strategy, it provides a relatively high standard reference policy for imitation learning.
To simulate variant levels of intelligent agents, we equip the privileged agents with both Long-Term Memory (LTM) and Short-Term Memory (STM). The STM retains only the most recent three pieces of observation, while the LTM retains information until the episode concludes. Only part of the STM is allowed to be recorded to LTM. We introduce a hyper-parameter, $p(\text{STM}\rightarrow\text{LTM})$, to govern the probability of transferring information from STM to LTM. At $p(\text{STM}\rightarrow\text{LTM})=0\%$, the agent operates without LTM, whereas $p(\text{STM}\rightarrow\text{LTM})=100\%$ allows for a fully functional LTM. For simplicity, a privileged agent with $\text{p}$ ( STM → LTM ) = p is written as preiviledged agent (p) for short.
@@
(b) The LTM of a privileged agent
Figure 5: A demonstration of the maze world tasks alongside privileged agents with smart navigation policy. The privileged agent is provided with limited access to the environment’s ground truth (b) under specific conditions. Only the areas within the agent’s line of sight are revealed, while the unsighted regions remain obscured. The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability. The privileged agent employs an exploration-then-exploitation strategy.